# CAST: Contrastive Adaptation and Distillation for Semi-Supervised Instance Segmentation

### v1

@@ Appendix A Introduction
Appendix B Related Work
@@
Appendix C Method
C.1 Overview
@@
Sec. C.2 formalizes our instance-aware pixel-wise contrastive loss, which is used in both Teacher Adaptation and Knowledge Transfer to enforce intra-instance cohesion and inter-instance separation; Sec. C.3 then details the three stages of the CAST pipeline.
C.2 Pixel-wise Contrastive Loss
@@
Negatives are sampled by our instance-aware sampler (§C.2), producing indices $\{(b^{\prime},q_{r})\}_{r=1}^{R}$ and corresponding similarities $s^{-}_{b,p,r}$.
@@
Assumption C.1 (Negative Sampling Guarantee).
When sampling a negative under our instance aware scheme, the probability it originates from a different instance is at least $p>0.5$, where $p$ can be estimated empirically (see Sec. D.3).
Proposition C.1 (Expected Margin Growth).
Under Assumption C.1, one gradient update on $\mathcal{L}_{\rm pxl}$ increases the expected inter-instance margin $\Delta_{\rm emp}$ by
@@
In practice, raising $\lambda_{\rm pxl}$ enhances margin growth but also increases training cost. If $\lambda_{\rm pxl}$ is too large, it can overemphasize inter-instance separation at the expense of intra-instance cohesion. We validate this effect in Sec. D.3 and provide a proof sketch in Appendix C.
C.3 CAST Framework
@@
Appendix D Experiments
D.1 Experimental Protocol
@@ D.2 Main Results
D.3 Empirical Validation
We validate Proposition C.1 by monitoring the false negative rate ($\mathrm{FNR}$), the fraction of sampled negatives that actually belong to the same instance, and the empirical margin
@@ D.3 Empirical Validation
Defining $p=1-\mathrm{FNR}$ as the success probability of sampling a true negative, Figure 3 shows: the empirical margin every 10 k iterations for $\lambda_{\rm pxl}\in\{0.01,0.05,0.1,0.2\}$ (left), the raw contrastive loss for $\lambda_{\rm pxl}=0.1$ (center), and the false negative rate for $\lambda_{\rm pxl}=0.1$ (right, dashed at $p=0.5$). Throughout training we observe $p>0.9$ and a linear increase of $\Delta_{\rm emp}$ with $\lambda_{\rm pxl}$, in agreement with Proposition C.1.
@@ Appendix E Conclusions
Appendix F Dataset Splits
@@ Appendix G Hyperparameters
Appendix H Proof Sketch of Proposition 3.1
@@
Remark H.1 (Why $\langle z^{+},z^{-}\rangle\approx 0$ holds).
@@
Appendix I More Training Details
@@
Appendix J Additional Ablations
J.1 Loss Variant: InfoNCE vs. Margin Hinge
@@
J.2 Ablation: Debias Score Formulation
@@
J.3 Ablation: Negative Sampling Scope

### v3

@@ Appendix 1 Introduction
Appendix 2 Related Work
@@
Appendix 3 Method
3.1 Overview
@@
Sec. 3.2 formalizes our instance-aware pixel-wise contrastive loss, which is used in both Teacher Adaptation and Knowledge Transfer to enforce intra-instance cohesion and inter-instance separation; Sec. 3.3 then details the three stages of the CAST pipeline.
3.2 Pixel-wise Contrastive Loss
@@
Negatives are sampled by our instance-aware sampler (§3.2), producing indices $\{(b^{\prime},q_{r})\}_{r=1}^{R}$ and corresponding similarities $s^{-}_{b,p,r}$.
@@
Assumption 3.1 (Negative Sampling Guarantee).
When sampling a negative under our instance aware scheme, the probability it originates from a different instance is at least $p>0.5$, where $p$ can be estimated empirically (see Sec. 4.3).
Proposition 3.1 (Expected Margin Growth).
Under Assumption 3.1, one gradient update on $\mathcal{L}_{\rm pxl}$ increases the expected inter-instance margin $\Delta_{\rm emp}$ by
@@
In practice, raising $\lambda_{\rm pxl}$ enhances margin growth but also increases training cost. If $\lambda_{\rm pxl}$ is too large, it can overemphasize inter-instance separation at the expense of intra-instance cohesion. We validate this effect in Sec. 4.3 and provide a proof sketch in Appendix C.
3.3 CAST Framework
@@
Appendix 4 Experiments
4.1 Experimental Protocol
@@ 4.2 Main Results
4.3 Empirical Validation
We validate Proposition 3.1 by monitoring the false negative rate ($\mathrm{FNR}$), the fraction of sampled negatives that actually belong to the same instance, and the empirical margin
@@ 4.3 Empirical Validation
Defining $p=1-\mathrm{FNR}$ as the success probability of sampling a true negative, Figure 3 shows: the empirical margin every 10 k iterations for $\lambda_{\rm pxl}\in\{0.01,0.05,0.1,0.2\}$ (left), the raw contrastive loss for $\lambda_{\rm pxl}=0.1$ (center), and the false negative rate for $\lambda_{\rm pxl}=0.1$ (right, dashed at $p=0.5$). Throughout training we observe $p>0.9$ and a linear increase of $\Delta_{\rm emp}$ with $\lambda_{\rm pxl}$, in agreement with Proposition 3.1.
@@ Appendix 5 Conclusions
Appendix 6 Dataset Splits
@@ Appendix 7 Hyperparameters
Appendix 8 Proof Sketch of Proposition 3.1
@@
Remark 8.1 (Why $\langle z^{+},z^{-}\rangle\approx 0$ holds).
@@
Appendix 9 More Training Details
@@
Appendix 10 Additional Ablations
10.1 Loss Variant: InfoNCE vs. Margin Hinge
@@ Appendix 10 Additional Ablations
10.2 Ablation: Debias Score Formulation
@@
10.3 Ablation: Negative Sampling Scope